Solve the formula for the specified variable. Write the
answer as an equation. NOTE: Maintain the case of the
given variables.
F=9/5C + 32 for C

Answer by stanbon(75887)

F=(9/5)C + 32 for C
--------------
(9/5)C = F-32
---
C = (5/9)*(F-32)
================
Cheers,
Stan H.